OpenStudy (anonymous):

Help ; 16 - (2y - 4) = 4y

OpenStudy (anonymous):

are you solving for y?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if so you would just need to do subtraction/addition \[20=6y\] and now some division \[\frac{ 20 }{ 6 }=y\] Simplify it to become \[\frac{ 10 }{ 3 }\]

OpenStudy (anonymous):

16-2y+4=4y 20=4y+2y 20=6y y=20/6 10/3

OpenStudy (anonymous):

Thank you